Arlington Corporation's financial statements (dollars and shares are in millions) are provided here.

Balance Sheets as of December 31
2019 2018
Assets
Cash and equivalents $  15,000 $  14,000
Accounts receivable 30,000 25,000
Inventories 27,640 22,000
  Total current assets $ 72,640 $ 61,000
Net plant and equipment 47,000 46,000
Total assets $119,640 $107,000
Liabilities and Equity
Accounts payable $ 10,200 $  8,500
Accruals 7,100 6,000
Notes payable 6,100 5,300
  Total current liabilities $ 23,400 $ 19,800
Long-term bonds 20,000 20,000
  Total liabilities $ 43,400 $ 39,800
Common stock (4,000 shares) 60,000 60,000
Retained earnings 16,240 7,200
  Common equity $ 76,240 $ 67,200
Total liabilities and equity $119,640 $107,000
Income Statement for Year Ending December 31, 2019
Sales $210,000
Operating costs excluding depreciation and amortization 160,000
EBITDA $ 50,000
Depreciation & amortization 8,000
EBIT $ 42,000
Interest 4,000
EBT $ 38,000
Taxes (25%) 9,500.00
Net income $ 28,500.00
Dividends paid 19,460.00

Enter your answers in millions. For example, an answer of $25,000,000,000 should be entered as 25,000. Round your answers to the nearest whole number, if necessary.

  1. What was net operating working capital for 2018 and 2019? Assume that all cash is excess cash; i.e., this cash is not needed for operating purposes.

    2018 $    million
    2019 $    million
  2. What was Arlington's 2019 free cash flow?

    $    million

  3. Construct Arlington's 2019 statement of stockholders' equity.

    Statement of Stockholders' Equity, 2019
    Common Stock Retained
    Earnings
    Total Stockholders'
    Equity
    Shares Amount
    Balances, 12/31/18   million $    million $    million $    million
    2019 Net Income   million
    Cash Dividends   million
    Addition to retained earnings   million
    Balances, 12/31/19   million $    million $    million $    million
  4. What was Arlington's 2019 EVA? Assume that its after-tax cost of capital is 10%. Round your answer to the nearest cent.

    $    million

  5. What was Arlington's MVA at year-end 2019? Assume that its stock price at December 31, 2019 was $25. Round your answer to the nearest cent.

    $    million
Answer #1

a.

2019 2018
Total current assets          72,640          61,000
Total current liabilities          23,400          19,800
Working capital          49,240          41,200
Less: cash          15,000          14,000
Net operating working capital          34,240          27,200

b. Free cash flow = Net operating profit after taxes (NOPAT) - Net investment in operating capital

= Net income (excluding depreciation & interest) - Net investment in operating capital

= 28,500+ 8,000+4,000 = 40,500 - 7,040 = 33,460

c.  Arlington's 2019 statement of stockholders' equity.

Statement of Stockholders' Equity, 2019
Common Stock Retained Total Stockholders'
Shares Amount Earnings Equity
Balances, 12/31/18 4000 60000 16240 76240
2019 Net Income 28500
Cash Dividends -19460
Addition to retained earnings
Balances, 12/31/19 85280

d. Arlington's 2019 EVA = NOPAT - WACC(Total assets - current liabilities)

= 33,460 - 10%(119,640-23,400) = 33,360 - 9,624

= 23,836